OpenStudy (sonyalee77):

A college student is looking at her monthly checking account records. On September 1, 2008, her checking account held a balance of $1,050. At the end of March 2009, her checking account held a balance of $800. What is the monthly rate of change for the student’s checking account over the period September 1, 2008 through March 31, 2009?

OpenStudy (sonyalee77):

\[A. -\frac{ 1,850 }{ 7} B. \frac{ -250 }{ 7} C.\frac{ 250 }{ 7 } D.\frac{ 1850 }{ 7 }\]

OpenStudy (sonyalee77):

@fujisaki @Euler271 @Nurali

OpenStudy (anonymous):

Sept 1 through the end of March is 7 complete months $800-$1050=$-250

OpenStudy (anonymous):